A right circular cone is inscribed in a cube having a diagonal which
measures 20 square root of 3 cm.

2. What is the volume of the cone?
1
Expert's answer
2020-12-02T01:44:42-0500

The diagonal "d" of the cube with the side "a" is


"d^2=a^2+a^2+a^2=3a^2"

"d=\sqrt{3}a"

Given "d=20\sqrt{3}\ cm."

Then


"a=\dfrac{d}{\sqrt{3}}=\dfrac{20\sqrt{3}}{\sqrt{3}}=20(cm)"

The radius "r" of the circle inscribed in the square with the side "a" is


"r=\dfrac{a}{2}"

Then we have the right circular cone inscribed in a cube with the side "a"

"radius=r=\dfrac{a}{2}, \ height=h=a"

The volume of the cone is


"V_{cone}=\dfrac{1}{3}\pi r^2h=\dfrac{1}{3}\pi (\dfrac{a}{2})^2 (a)=\dfrac{\pi a^3}{12}"

"V_{cone}=\dfrac{\pi(20)^3}{12}=\dfrac{100\pi}{3}(cm^3)"

"\approx104.720(cm^3)"

The volume of the cone is

"\dfrac{100\pi}{3}\ cm^3\approx 104.720 \ cm^3."